OpenStudy (anonymous):

A bank offers two interest account plans. Plan A gives you 6% interest compound annually. Plan B gives you 13% annual simple interest. You plan to invest $2,000 for the next four years. Which account earns you the most interest (in dollars) after four years? How much will you have earned? A.) Plan A; $524.95 B.) Plan B; $524.95 C.) Plan B; $1,040.00 D.) Plan A; $1,040.00

hero (hero):

Annual compound formula \[A = P(1 + r)^n\] P = start up amount n = years r = rate in decimal form Simple interest I = PRT I = interest P = principle R = rate (in decimal form) T = time in years Plug in the values for both, then calculate